ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 204, 2014

Aistleitner, C. Quantitative uniform distribution results for geometric progressions

155

Bae, Y. J. On the growth sequence for symplectically hyperbolic manifolds

417

Ballester-Bolinche, A., R. Esteban-Romero and L. M. Ezquerro On the p-length of some finite p-soluble groups

359

Caviglia, G., A. Constantinescu and M. Varbaro On a conjecture by Kalai

469

Chen, Q. and D. Wang A Maschke type theorem for weak group entwined modules and applications

329

Chen, S., A. Rasila and X. Wang Radial growth, Lipschitz and Dirichlet spaces on solutions to the non-homogenous Yukawa equation

261

Chernikov, A. and M. Hils Valued difference fields and NTP_{2}  299

Constantinescu, A. See G. Caviglia Esteban-Romero, R. See A. Ballester-Bolinche Ezquerro, L. M. See A. Ballester-Bolinche Gishboliner, L. and A. Shapira Deterministic vs non-deterministic

graph property testing

397

Guo, X. See G. Liu Hils, M. See A. Chernikov Hric, R. and T. Jäger A construction of almost automorphic min-

imal sets

373

Jäger, T. See R. Hric Jantzen, C. and B. Liu The generic dual of p-adic split SO_{2n} and

local langlands parameters

199

Jiang, Q. See C. Shao Liu, B. See C. Jantzen Liu, G. and X. Guo Harish-Chandra modules over generalized

Heisenberg–Virasoro algebras

447

Martín-Morales, J. Embedded Q-resolutions for Yomdin–Lê surface singularities

97

Montalbán, A. and R. A. Shore The limits of determinacy in second order arithmetic: consistency and complexity strength 477 Rasila, A. See S. Chen Shao, C. and Q. Jiang An extension of a theorem of Alan Camina

on conjugacy class sizes

145

Shapira, A. See L. Gishboliner

Shelah, S. Strongly dependent theories  1

Shore, R. A. See A. Montalbán Varbaro, M. See G. Caviglia Wang, D. See Q. Chen Wang, X. See S. Chen Yang, C. F. An inverse problem for a differential pencil using nodal

points as data

431

Yang, C. F. Solutions to open problems of Yang concerning inverse nodal problems

283

Zorin-Kranich, P. Return times theorem for amenable groups  85